OpenStudy (anonymous):

Simplify −24x^6y^7/18x^−3y^9

OpenStudy (apoorvk):

\[\large a^-n = \frac{1}{a^{n}}\] \[\large a^m * a^n = a^{m+n}\] \[\large \frac {a^m}{a^n} = a^{m-n}\] apply this for every term and 'x' and 'y', and simplify